Could you please help me answer this question using quadratic equations:
2x(4x-5)=3
Thanks in advance.

Answer by scott8148(6628)

distributing and subtracting 3 __ 8x^2-10x-3=0

factoring (4x+1)(2x-3)=0 __ 4x+1=0 and 2x-3=0

so, x=-1/4 and x=3/2